# The new observations about the parameter-dependent Schrödinger-Poisson system

### v1

@@ 4.2. Critical groups at zero
In this section, we will use the following proposition to compute the critical groups of $I_{\epsilon}$ at zero.
@@
Under assumptions ${\bf(V_{1})}$, ${\bf(V_{2})}$, ${\bf(f)}$ and ${\bf(f^{\prime})}$. Then the functional $I_{\epsilon}$ has a local linking at zero with respect to decomposition $H^{1}_{V}(\mathbb{R}^{3})=H^{1}_{V}(\mathbb{R}^{3})^{-}\oplus H^{1}_{V}(\mathbb{R}^{3})^{+}$.
@@
Recall (2.2), for any $\delta>0$ and $\epsilon\in(0,\frac{\delta^{2}_{0}}{4})$, there is a constant $C_{\delta}$ such that
$0\leq\tilde{F}(t)\leq\delta t^{2}+C_{\delta}|t|^{3},$
where $C_{\delta}$ is a constant not depending on $\epsilon$. Hence, as $\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}\to 0$, we get
$\int_{\mathbb{R}^{3}}\tilde{F}(u)\text{d}x\leq o(\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}).$
Thus for $u\in H^{1}_{V}(\mathbb{R}^{3})^{+}\setminus\{\bf 0\}$ and $\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}$ small enough, by (4.3) we obtain
$I_{\epsilon}(u)$ $=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x$
$\geq\frac{\eta}{4}\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\frac{C}{4}\|u\|^{4}_{H^{1}_{V}(\mathbb{R}^{3})}+o(\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})$
$\geq\frac{\eta}{8}\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}.$
Then for $r>0$ small enough, we obtain
(4.5) $I_{\epsilon}(u)>0\ \mbox{for}\ u\in(H^{1}_{V}(\mathbb{R}^{3})^{+}\setminus\{{\bf 0}\})\cap B_{r}.$
On the other hand, for $u\in H^{1}_{V}(\mathbb{R}^{3})^{-}\cap B_{r}$, it is easy to see that
$I_{\epsilon}(u)$ $=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x$
$\leq-\eta\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x$
@@
From this and (4.5), it implies $I_{\epsilon}$ has a local linking property at zero. ∎
@@ 4.3. Critical groups at infinity
If assumptions ${\bf(V_{1})}$, ${\bf(V_{2})}$, ${\bf(f)}$ and ${\bf(f^{\prime})}$ hold. Then $C_{q}(I_{\epsilon},\infty)\cong 0$ with $q\in\mathbb{N}=\{0,1,2,\cdots\}$.
@@
(4.6) $I_{\epsilon}(sh)\to-\infty\ \mbox{as}\ s\to+\infty\ \mbox{for any}\ h\in S^{\infty}.$
Similar with Lemma 4, from $\tilde{F}(t)\geq 0$ we deduce
$\begin{split}I_{\epsilon}(sh)&=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla(sh)|^{2}+\tilde{V}(x)(sh)^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{(sh)}(sh)^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon sh)\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}(sh)^{2}\text{d}x\\ &\leq s^{2}\left(\frac{1}{2}-\frac{s^{2}}{4}\int_{\mathbb{R}^{3}}\phi_{h}h^{2}\text{d}x\right)\to-\infty.\end{split}$
@@
Claim: There is a constant $A>0$ such that if $I_{\epsilon}(u)\leq-A$, then
$\frac{\mbox{d}}{\mbox{d}t}|_{t=1}I_{\epsilon}(tu)<0.$
@@
(4.7) $I_{\epsilon}(u_{n})\leq-n\ \mbox{and}\ \langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle=\frac{\mbox{d}}{\mbox{d}t}|_{t=1}I_{\epsilon}(tu_{n})\geq 0.$
From (4.7), similar with (3.2), we have
(4.8) $\begin{split}0&\geq I_{\epsilon}(u_{n})-\frac{1}{3}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &=\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{n})\text{d}x\\ &\quad\quad-\frac{1}{3\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\left(1+\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(\delta|u_{n}|^{2}+\epsilon C_{\delta}|u_{n}|^{3}\right)\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\frac{1}{8}\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x,\end{split}$
where $\delta$ denoted by $\frac{1}{64}$ and $\epsilon>0$ small.
Notice that if $\{u_{n}\}_{n=1}^{\infty}$ is a bounded sequence in $H^{1}_{V}(\mathbb{R}^{3})$, then $\{I_{\epsilon}(u_{n})\}$ is also bounded. Then by $I_{\epsilon}(u_{n})\leq-n$, it implies
$\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\widetilde{V}(x)|u_{n}|^{2}\right)\text{d}x\to+\infty\ \mbox{as}\ n\to+\infty.$
@@
$\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mv^{2}_{n}\text{d}x\geq\frac{1}{8}.$
@@
By the definition of $\tilde{F}(t)$, conditions $({\bf f})$ and $({\bf f^{\prime}})$, for any $\epsilon\in(0,\frac{\delta_{0}}{4})$, it follows that
(4.9) $\tilde{F}^{\prime}(t)t\geq 0.$
@@
$\begin{split}&\indent\int_{\mathbb{R}^{3}}\left(\phi_{u_{n}}u^{2}_{n}+\frac{1}{\epsilon}\tilde{F}^{\prime}(\epsilon u_{n})u_{n}+mu_{n}^{2}\right)\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{-2}\text{d}x\\ &=\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{u_{n}^{2}}+\frac{\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}}{\epsilon^{2}u_{n}^{2}}+\frac{mu_{n}^{2}}{u_{n}^{2}}\right)v^{2}_{n}\text{d}x\\ &\geq\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\to+\infty\end{split}$
@@
$\begin{split}0&\leq\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &=\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}\text{d}x-m\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x\\ &\leq\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}\left(\frac{1}{2}-\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\right)\to$-$\infty\end{split}$
@@
$I_{\epsilon}(T(u)u)=-B\ \ \text{for every }u\in S^{\infty}.$
@@
with $\eta(0,u)=u$, $\eta(1,u)\in I_{\epsilon}^{-B}$. Thus
$C_{q}(I_{\epsilon},\infty)=H^{q}(H^{1}_{V}(\mathbb{R}^{3}),I_{\epsilon}^{-B})\cong H^{q}(H^{1}_{V}(\mathbb{R}^{3}),H^{1}_{V}(\mathbb{R}^{3})\setminus B_{1})\cong 0,\ \mbox{for all}\ q\in\mathbb{N}.$
@@
[…]

### v3

@@ 4.2. Critical groups at zero
In this section, we will use the following proposition to compute the critical groups of $I_{\epsilon,R}$ at zero.
@@
Under assumptions ${\bf(V_{1})}$, ${\bf(V_{2})}$, ${\bf(f)}$ and ${\bf(f^{\prime})}$. For any $R>1$ fixed, there exists $\epsilon^{\prime\prime}(R)>0$ such that for any $\epsilon\in(0,\epsilon^{\prime\prime}(R))$ then the functional $I_{\epsilon,R}$ has a local linking at zero with respect to decomposition $H^{1}_{V}(\mathbb{R}^{3})=H^{1}_{V}(\mathbb{R}^{3})^{-}\oplus H^{1}_{V}(\mathbb{R}^{3})^{+}$.
@@
Follows from the proof of Lemma 5. There exists $\epsilon^{\prime\prime}(R)>0$, such that for any $\epsilon\in(0,\epsilon^{\prime\prime}(R))$ and $u\in H^{1}_{V}(\mathbb{R}^{3})^{+}\setminus\{\bf 0\}$ with $\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}=r$, taking $r>0$ small enough, by (4.2) we obtain
$I_{\epsilon,R}(u)$ $=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x$
$\geq\frac{\eta}{4}\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\frac{C}{4}\|u\|^{4}_{H^{1}_{V}(\mathbb{R}^{3})}$
$\quad-\left(\delta\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}u^{2}\text{d}x+C_{\delta,R}\epsilon^{p-2}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left|u\right|^{p}\text{d}x\right)$
[…]
$\geq\frac{\eta}{16}\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}=\frac{\eta r^{2}}{16}>0.$
Then we obtain
(4.4) $I_{\epsilon,R}(u)>0\ \mbox{for}\ u\in(H^{1}_{V}(\mathbb{R}^{3})^{+}\setminus\{{\bf 0}\})\cap B_{r}.$
On the other hand, for $u\in H^{1}_{V}(\mathbb{R}^{3})^{-}\cap B_{r}$ and $\epsilon\in(0,\frac{\delta_{0}}{R})$, it is easy to see that
$I_{\epsilon,R}(u)$ $=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x$
$\leq-\eta\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u\right|\right)u\right)\text{d}x$
@@
From this and (4.4), it implies $I_{\epsilon,R}$ has a local linking property at zero. ∎
@@ 4.3. Critical groups at infinity
If assumptions ${\bf(V_{1})}$, ${\bf(V_{2})}$, ${\bf(f)}$ and ${\bf(f^{\prime})}$ hold. For any $R>1$ fixed, there exists a constant $\epsilon(R)>0$ such that $C_{q}(I_{\epsilon,R},\infty)\cong 0$ with $q\in\mathbb{N}=\{0,1,2,\cdots\}$.
@@
(4.5) $I_{\epsilon,R}(sh)\to-\infty\ \mbox{as}\ s\to+\infty\ \mbox{for any}\ h\in S^{\infty}.$
Similar with Lemma 6, for any $\epsilon\in(0,\frac{\delta_{0}}{R})$, from $F\left(\epsilon\eta_{R}\left(\left|t\right|\right)t\right)\geq 0$ we deduce
$\begin{split}I_{\epsilon,R}(sh)&=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla(sh)|^{2}+\tilde{V}(x)(sh)^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{(sh)}(sh)^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|sh\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|sh\right|\right)sh\right)\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}(sh)^{2}\text{d}x\\ &\leq s^{2}\left(\frac{1}{2}-\frac{s^{2}}{4}\int_{\mathbb{R}^{3}}\phi_{h}h^{2}\text{d}x\right)\to-\infty,~~\text{as}~s\to+\infty.\end{split}$
@@
Claim: There is a constant $A>0$ such that if $I_{\epsilon,R}(u)\leq-A$, then
$\frac{\mbox{d}}{\mbox{d}t}|_{t=1}I_{\epsilon,R}(tu)<0.$
@@
(4.6) $I_{\epsilon,R}(u_{n})\leq-n\ \mbox{and}\ \langle I_{\epsilon,R}^{\prime}(u_{n}),u_{n}\rangle=\frac{\mbox{d}}{\mbox{d}t}|_{t=1}I_{\epsilon,R}(tu_{n})\geq 0.$
Notice that if $\{u_{n}\}_{n=1}^{\infty}$ is a bounded sequence in $H^{1}_{V}(\mathbb{R}^{3})$, then $\{I_{\epsilon,R}(u_{n})\}$ is also bounded. Then by $I_{\epsilon,R}(u_{n})\leq-n$, it implies
$\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\widetilde{V}(x)u_{n}^{2}\right)\text{d}x\to+\infty\ \mbox{as}\ n\to+\infty.$
From (4.6), similar with Lemma 5, we have
(4.7) $\begin{split}0&\geq I_{\epsilon,R}(u_{n})-\frac{1}{3}\langle I_{\epsilon,R}^{\prime}(u_{n}),u_{n}\rangle\\ &=\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\quad\quad-\frac{2}{3\epsilon^{2}}\eta_{R}^{\prime}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2})\text{d}x\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\text{d}x\\ &\quad\quad-\frac{1}{3\epsilon}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left(f\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\left(\eta_{R}\left(\left|u_{n}\right|\right)u_{n}+\eta_{R}^{\prime}\left(\left|u_{n}\right|\right)u_{n}^{2}\right)\right)\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\frac{4}{3}\left(\frac{1}{R}+C_{R}\epsilon^{p-2}\right)-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x-\frac{4}{3},\end{split}$
where $\epsilon\in(0,\epsilon(R))$.
@@
$\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mv^{2}_{n}\text{d}x\geq\frac{1}{6}.$
@@
It follows from assumptions $({\bf f})$ and $({\bf f^{\prime}})$ that for $\epsilon\in(0,\frac{\delta_{0}}{R})$,
(4.8) $F\left(\epsilon\eta_{R}\left(\left|t\right|\right)t\right)\geq 0\ ~\text{for all}~t\in\mathbb{R}.$
@@
$\begin{split}&\indent\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4}+\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)+\frac{1}{2}mu_{n}^{2}\right)\rho^{-2}_{n}\text{d}x\\ &\geq\int_{\mathbb{R}^{3}}\frac{\phi_{u_{n}}u^{2}_{n}}{4u_{n}^{2}}v^{2}_{n}\text{d}x\\ &\geq\frac{1}{4}\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\to+\infty,\end{split}$
@@
$\begin{split}0&\leq\langle I_{\epsilon,R}^{\prime}(u_{n}),u_{n}\rangle\\ &=\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{n}\right|\right)u_{n}\right)\text{d}x-m\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x\\ &\leq\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}\left(1-\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\right)\to-\infty,\end{split}$
@@
$I_{\epsilon,R}(T(u)u)=-B\ \ \text{for every }u\in S^{\infty}.$
@@
with $\eta(0,u)=u$, $\eta(1,u)\in I_{\epsilon,R}^{-B}$. Thus
$C_{q}(I_{\epsilon,R},\infty)=H^{q}(H^{1}_{V}(\mathbb{R}^{3}),I_{\epsilon,R}^{-B})\cong H^{q}(H^{1}_{V}(\mathbb{R}^{3}),H^{1}_{V}(\mathbb{R}^{3})\setminus B_{1})\cong 0,\ \mbox{for all}\ q\in\mathbb{N}.$
@@
Proof of Theorem 2 It follows from Lemma 6 that $I_{\epsilon,R}$ satisfies the Cerami conditions. By Lemma 7, $I_{\epsilon,R}$ has a local linking at zero with respect to the decomposition $H^{1}_{V}(\mathbb{R}^{3})=H^{1}_{V}(\mathbb{R}^{3})^{-}\oplus H^{1}_{V}(\mathbb{R}^{3})^{+}$, thus from Proposition 2, for $d=\dim H^{1}_{V}(\mathbb{R}^{3})^{-}$, it implies $C_{d}(I_{\epsilon},0)\neq 0.$ On the other hand, due to Lemma 8, we have for all $q\in\mathbb{N}$, $C_{q}(I_{\epsilon,R},\infty)=0$. Thus, by Proposition 1, for any $R>1$, there exists $\epsilon^{\ast}(R)=\min\{\epsilon(R),\epsilon^{\prime\prime}(R)\}$ for any $\epsilon\in(0,\epsilon^{\ast}(R))$ such that $I_{\epsilon,R}$ has a nontrivial critical point $u_{\epsilon,R}$.
By the proof of Lemma 5, for any $R>1$ fixed, one has
$\begin{split}&\lim\limits_{\epsilon\to 0^{+}}\left(I_{\epsilon,R}(u_{\epsilon,R})-I_{0}(u_{\epsilon,R})\right)\\ =&\lim\limits_{\epsilon\to 0^{+}}\frac{1}{\epsilon^{2}}\eta_{R}\left(\|u_{\epsilon,R}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{\epsilon,R}\right|\right)u_{\epsilon,R}\right)\text{d}x\\ =&0.\end{split}$
Similarly, for any $v\in H^{1}_{V}(\mathbb{R}^{3})$, it implies
$\begin{split}&\lim\limits_{\epsilon\to 0^{+}}\langle I_{\epsilon,R}^{\prime}(u_{\epsilon,R})-I_{0}^{\prime}(u_{\epsilon,R}),u_{\epsilon,R}+v\rangle\\ =&\lim\limits_{\epsilon\to 0^{+}}\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}\left(\|u_{\epsilon,R}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}(|\nabla u_{\epsilon,R}|^{2}+\tilde{V}(x)u^{2}_{\epsilon,R})\text{d}x\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{\epsilon,R}\right|\right)u_{\epsilon,R}\right)\text{d}x\\ &+\lim\limits_{\epsilon\to 0^{+}}\frac{1}{\epsilon}\eta_{R}\left(\|u_{\epsilon,R}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left(f\left(\epsilon\eta_{R}\left(\left|u_{\epsilon,R}\right|\right)u_{\epsilon,R}\right)\left(\eta_{R}\left(\left|u_{\epsilon,R}\right|\right)u_{\epsilon,R}+\eta_{R}^{\prime}\left(\left|u_{\epsilon,R}\right|\right)|u_{\epsilon,R}|u_{\epsilon,R}\right)\right)\text{d}x\\ =&\lim\limits_{\epsilon\to 0^{+}}\frac{2}{\epsilon^{2}}\eta_{R}^{\prime}\left(\|u_{\epsilon,R}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}(\nabla u_{\epsilon,R}\nabla v+\tilde{V}(x)u_{\epsilon,R}v)\text{d}x\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{R}\left(\left|u_{\epsilon,R}\right|\right)u_{\epsilon,R}\right)\text{d}x\\ &+\lim\limits_{\epsilon\to 0^{+}}\frac{1}{\epsilon}\eta_{R}\left(\|u_{\epsilon,R}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}\left(f\left(\epsilon\eta_{R}\left(\left|u_{\epsilon,R}\right|\right)u_{\epsilon,R}\right)\left(\eta_{R}\left(\left|u_{\epsilon,R}\right|\right)v+\eta_{R}^{\prime}\left(\left|u_{\epsilon,R}\right|\right)|u_{\epsilon,R}|v\right)\right)\text{d}x\\ =&0.\end{split}$
Thus when $\epsilon\to 0^{+}$, $\{u_{\epsilon,R}\}$ is a Cerami sequence of $I_{0}$. Then by the Cerami condition of $I_{0}$, we have $u_{\epsilon,R}\to u_{0}$ strongly in $H^{1}_{V}(\mathbb{R}^{3})$ as $\epsilon\to 0^{+}$, moreover $\{u_{\epsilon,R}\}$ is uniformly bounded in $H^{1}_{V}(\mathbb{R}^{3})$. Similar with Lemma 4, we can obtain the $L^{\infty}$ norm of $u_{\epsilon,R}$ is uniformly bounded about $\epsilon$. This means that there exists a positive constant $M$ independent of $\epsilon$, such that for any $\epsilon\in(0,\epsilon^{\ast}(R))$ one has
$\max\{\|u_{\epsilon,R}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})},\|u_{\epsilon,R}\|_{L^{\infty}(\mathbb{R}^{3})}\}\leq M.$
Finally, let $R=M+1$, for any $\epsilon\in(0,\epsilon^{\ast}(M+1))$, then
$\begin{split}I_{\epsilon,M+1}\left(u_{\epsilon,M+1}\right)&=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u_{\epsilon,M+1}|^{2}+V(x)u_{\epsilon,M+1}^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u_{\epsilon,M+1}}u_{\epsilon,M+1}^{2}\text{d}x\\ &\quad-\frac{1}{\epsilon^{2}}\eta_{M+1}\left(\|u_{\epsilon,M+1}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}\right)\int_{\mathbb{R}^{3}}F\left(\epsilon\eta_{M+1}\left(\left|u_{\epsilon,M+1}\right|\right)u_{\epsilon,M+1}\right)\text{d}x\\ &=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u_{\epsilon,M+1}|^{2}+V(x)u_{\epsilon,M+1}^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u_{\epsilon,M+1}}u_{\epsilon,M+1}^{2}\text{d}x\\ &\quad-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}F\left(\epsilon u_{\epsilon,M+1}\right)\text{d}x\end{split}$
has a nontrivial critical point $u_{\epsilon,M+1}$. Hence, denote $\epsilon^{**}=\epsilon^{\ast}(M+1)$, then for any $\epsilon\in(0,\epsilon^{**})$, $u_{\epsilon,M+1}$ is also the solution of (1.2). Consequently, $\epsilon u_{\epsilon,M+1}$ is the solution of (1.1) and $\lambda^{**}=\frac{1}{(\epsilon^{**})^{2}}$. Thus Theorem 2 has been proved.